Write the equation of a line passing through the points (3,-2) and (-4,-1)

Answer by lwsshak3(11628)

Write the equation of a line passing through the points (3,-2) and (-4,-1)
***
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
slope=∆y/∆x=(-1-(-2))/(-4-3)=1/-7=-1/7
equation: y=-x/7+b
solving for b using coordinates of one of given points (3,-2)
-2=-3/7+b
b=-2+3/7=-11/7
Equation of given line: y=-x/7-11/7